OpenStudy (anonymous):

WILL GIVE MEDALS PLEASE HELP Shiloh is 7 years older than Courtney. Shiloh's age is 13 years less than two times Courtney's age. The system below models the relationship between Shiloh's age (s) and Courtney's age (c): s = c + 7 s = 2c − 13 Which of the following methods is correct to find Shiloh's and Courtney's age? Solve s + 7 = 2s − 13 to find the value of s. Solve c + 7 = 2c − 13 to find the value of c. Write the points where the graphs of the equations intersect the x-axis. Write the points where the graphs of the equations intersect the y-axis.

OpenStudy (anonymous):

#2 Solve c + 7 = 2c − 13 to find the value of c.

OpenStudy (anonymous):

can you explain

OpenStudy (skullpatrol):

s=s

OpenStudy (skullpatrol):

s = c + 7 s = 2c − 13

OpenStudy (anonymous):

If s = c + 7 s = 2c − 13 then one can equate the RHS of each equation to the other.